1. A boy throws a ball horizontally from the top of a building. The initial speed of the ball is 20 m/s. Find the vertical and horizontal components of its velocity. 2. A regulation volleyball court is 18.0 m long and a regulation volleyball net is 2.43 m high. Alyssa Valdez has to make a serve so she strikes the ball a height h = 1.73 m directly above the back line, and the ball's initial velocity makes an angle 50 degrees with respect to the ground. Will she be able to make one good serve? Prove your answer quantitatively. 3. An object is whirled at the end of a string in a horizontal circle. If the period of one revolution is 1.18 s, and the length of the sting is 0.93 m, find the centripetal acceleration of the object.
1
Expert's answer
2021-10-25T09:56:20-0400

1)

"v_x=v,"

"v_y=gt,"

2)

"L_{max}=\frac{v^2\sin 2\alpha}g=40.2~m>d=18~m,"

"h_{max}=\frac{v^2\sin^2\alpha}{2g}=11.9~m>l-h=0 .7~m,"

she will not be able to make one good serve,

3)

"a=\omega^2r=\frac{4\pi^2r}{T^2}=26.3~\frac m{s^2}."